# Gravity-Informed Deep Learning Framework for Predicting Ship Traffic Flow and Invasion Risk of Non-Indigenous Species via Ballast Water Discharge

### v1

@@ Abstract
Aquatic invasive species are a significant threat to biodiversity and ecosystems worldwide. Due to the fast growth of global trade and transportation networks, non-native species have been introduced and spread unintentionally in new environments. These invasive species can damage ecosystems by disrupting the balance, out-competing native species, and causing significant economic losses to agriculture, forestry, and fisheries. Therefore, there is a pressing need for risk assessment and management techniques to mitigate the impact of these invasions. This study aims to develop a new physics-inspired model to forecast maritime shipping traffic and thus inform risk assessment of invasive species spread through global transportation networks. Inspired by the gravity model for international trades, our model considers various factors that influence the likelihood and impact of vessel activities, such as shipping flux density, distance between ports, trade flow, and centrality measures of transportation hubs. Additionally, by analyzing the risk network of invasive species, we provide a comprehensive framework for assessing the invasion threat level given a pair of origin and destination. Accordingly, in this paper, we introduce transformers to gravity models to rebuild the short- and long-term dependencies that make the risk analysis feasible. Thus, we introduce a physics-inspired framework that achieves an 89% segmentation accuracy for existing and non-existing trajectories and an 84.8% accuracy for the number of vessels flowing between key port areas, representing more than 10% improvement over the traditional deep-gravity model. Along these lines, this research contributes to a better understanding of invasive species risk assessment. It allows policymakers, conservationists, and stakeholders to prioritize management actions by identifying high-risk invasion pathways. Besides, our model is versatile and can include new data sources, making it suitable for assessing species invasion risks in a changing global landscape.
@@ Introduction
Globalization has rapidly increased marine shipping activities in the last decades. According to a statistics report, container shipping has increased by 24 times in tonnage from 1980 to 2020 [1]. During this time, the environmental pollution caused by introducing Non-Indigenous Species (NIS) through shipping activities has been a subject of study regarding marine preservation. Ballast water, used to keep vessels in balance during travels, is listed as a major source of NIS pollution [2]. The introduction of NIS into different ecological regions due to ballast water discharge has been shown to cause significant damage to the local ecosystem. It poses a severe threat to the biodiversity of affected areas. In response to biological invasion issues, many studies about Ballast Water Risk Assessment (BWRA) and tools have been developed over the last decades to estimate the risk levels of carrying NIS in the ballast water tanks [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. These works and tools rely on ship self-reports, such as made available by the National Ballast Information Clearinghouse (NBIC) [14]. Ballast water reporting forms provide information on the water source and destination areas, allowing for an assessment of the environmental similarity between source and target locations of a vessel voyage, which is considered a significant indicator of invasion risk level [2]. However, the acquisition of ballast reports is limited at the global scale due to the various policies across different countries. Additionally, BWRA lacks alternative sources that incorporate comprehensive shipping information.
@@
Recent research has revealed a strong correlation between the introduction of non-indigenous species and the movement of ships through shipping networks. These studies [12, 15] have utilized data from the Automatic Identification System (AIS), a sophisticated location tracking system on ships that allows them to share their positions in real-time [16]. This technology allows researchers to track individual/collective ships [17, 18, 19, 20], predict larger-scale shipping activities [21], and assess the risk of introducing NIS through ballast water [12]. AIS data has emerged as a promising source of information for studying bioinvasions in marine ecosystems. These studies analyze mobility flow by using Origin-Destination (OD) models that combine physics with statistical mechanics. The gravity model, inspired by Newton’s Law of Universal Gravitation, measures the attractive force between two objects based on their masses and the distance between them [22]. The gravity-inspired OD models were introduced in early human mobility and migration studies [23, 24]. They rely on information about population size and distances between origins and destinations as features.
The gravity theory permeates many areas of study that go beyond mobility and migration, such as the spreading of epidemics [25, 26], commercial trading [27], communication [28] and shipping networks [15, 29] modeling. Although the gravity model has been widely used to model real-world problems, recent studies have shown that it may not be sufficient for capturing complex patterns in various scenarios [30]. Relying solely on mass and distance as the critical factors of the model could lead to failures in accurately representing patterns [31]. Nonetheless, the gravity model has been prevalent for many years and remains a popular tool for modeling various phenomena. Beyond gravity models, radiation absorption is another physics-inspired OD model to study mobility patterns [32, 33, 34]. Unlike gravity models, which draw inspiration from gravitational forces, radiation models are based on principles seen in radiation and absorption processes, also from physics. While gravity models contain adjustable variables that may be difficult to define, the radiation-absorption model simplifies this by emphasizing distance as the primary feature while considering the population density. The well-known application of the radiation model is to predict human movement toward locations influenced by employment indicators [32]. Further studies used field theory for abstracting vector fields of daily commuting flows [35], while others translated field theory-based mobility to deep learning models for achieving better interpretation of spatiotemporal features in mobility patterns [36].
@@
The deep learning techniques discussed above in the mobility domain are a subset of the machine learning realm, and they have become increasingly popular in various applications due to their ability to recognize patterns by fine-tuning multiple parameters. These techniques have been used to forecast vessel trajectories in the ocean, predict patient trajectories in hospitals, track the spread of epidemics, and many other applications [37, 38, 39, 40]. Understandably, coupling machine or deep learning capabilities for pattern recognition with physics-inspired OD models can imply a higher proficiency in capturing and predicting complex scenarios. For instance, a study in human mobility used deep learning methods with the gravity model [41]. The resulting composite model called the Deep Gravity Model, has expanded the standard feature set of conventional gravity models, which typically incorporated population size and distance between OD locations, to include a variety of parameters characterizing the origins and destinations such as land-use patterns and the presence of retail and healthcare amenities. A similar study proposed the use of machine learning models augmented with urban indicators to forecast the flow of commuters during worker migrations [42], which included data about labor, safety, and quality of life and work in the cities using censuses-collected data. The enriched feature set enabled both models to discern and capture the intricacies of human mobility across wider information spaces. Specifically, deep learning techniques have inherent learning mechanics characterized by feed-forward and feedback-connectivity. These mechanics allow for adaptive weight assignment to individual variables. Therefore, this blend of deep learning and traditional OD models sets a promising premise for its application in maritime traffic mobility networks, as it allows intricate exploration of patterns at a granular level required for shipping network analysis.
@@
Along these lines, this paper focuses on predicting the OD fluxes of marine vessels to gain insight into global shipping behavior and their role in preventing NIS cases through BWRA. Ballast water is essential in maritime operations, particularly for container ships. As illustrated in Figure 1, while ballast water helps stabilize the ship’s load, local species can enter the ballast tanks and travel long distances when the water is taken in and discharged. Therefore, forecasting shipping patterns is critical to understanding the risk of spreading NIS. To this end, we propose a gravity-informed model where the shipping fluxes are considered as “mass”, and the vessel traffic flows are inversely proportional to the distances traveled by ships. We have enriched our model with relevant features from shipping activities, including bilateral trade data between countries [43] and graph metrics extracted from the global shipping network. Our deep learning model, known as the Transformer Gravity, relies on the transformer architecture [44] and is proficient in capturing local and global data dependencies through self-attention mechanisms. This mechanism enables the model to weigh the importance of different parts of the input sequence, assigning varying degrees of attention depending on the relevance of each input part during generating the output. As a result, our model can discern and incorporate short- and long-term dependencies in vessel traffic flows, making it more sensitive to the complex and dynamic patterns in maritime vessel movements.
@@
As part of our proposed framework, we have employed a machine learning classifier that proceeds the flow estimation process and filters out non-existing edges on a link-prediction binary classification task, a two-stage modeling process. This allows only highly probable flows, based on prior knowledge, to be fed into the gravity-based models, where the final flow estimations take place with the aid of the gravity-based model. In this sense, we have conducted experiments using the (a) Transformer Gravity, (b) Deep Gravity, and (c) Shallow Gravity Models. We utilized regression models in machine learning for performance comparison, and the same approach was used regarding the binary classification task. Our results demonstrate that the Transformer Gravity model significantly outperforms all the other approaches as it achieves an average Common Part of Commuters (CPC) of 85.3%, representing an improvement higher than 10% in the model output certainty in contrast to the Deep Gravity counterpart and over 20% improvement compared to the Shallow Gravity and other regression models. The results we have obtained are due not only to the proposed model but also to the proposed pipeline. We have significantly improved performance by incorporating prior knowledge about potential destinations and using the attention mechanism and the traditional gravity-informed model for mobility flow estimation.
Overall, we provide a consistent advancement in the gravity-informed flow estimation models that paved the way for ballast water risk assessment and management by enabling the forecasting of vessel mobility flows. That was possible mainly due to advancements in the model pipeline and architecture, which strongly rely upon key ocean and vessel mobility domain features. Specifically, data from global economic trades between countries and graph centrality metrics from port networks significantly contribute to achieving state-of-the-art results. This emphasizes the value of integrating shipping network analysis and trade information into vessel mobility flow predictions. Although gravity-informed models have limitations in capturing temporal dynamics [30], the annual patterns in ship traffic flows predicted by the Transformer Gravity model can provide references for shipping intensity in ballast water risk assessment. In addition, due to the versatility of mobility data, we believe that our framework and model can be utilized in various fields of research such as human mobility, urban transportation, and epidemic modeling.
@@ Results
At the beginning of this study, we constructed a directed and weighted international shipping network based on global port visits data in the same way as in [45] from 2017 to 2019. Ports and shipping connections were represented as nodes and edges in the shipping network, and the World Port Index (WPI) [46] was also used for port identification. Figure 2 provides an overview of global shipping connections in the three years from 2017 to 2019 and the distribution of ports participating in these shipping activities.
@@
Notice that, our study was conducted during a period of significant events (e.g., war, global pandemic, and economic sanctions) that could have influenced the underlying shipping data. However, all results we discuss use the same methodology, ensuring that any external factors would have had a homogeneous impact on the comparative performance of the models, maintaining the integrity of our analysis and results.
Figure 2: Global shipping network joining data from 2017 to 2019.
Global Shipping Network
The global shipping network can be described as $G=(V,E,W)$, where $V$ represents the set of ports, $E$ is the set of shipping routes connecting pairs of ports, and $W$ is the collection of edge weights. In this context, each weight in $W$ corresponds to the number of individual trips $T_{ij}$ between port $i$ and port $j$:
$W=\{w_{ij}:w_{ij}=\sum_{t}T_{ij}(t)\,\forall(i,j)\in E\}$
Figure 3: Pipeline for analyzing and predicting links in the global shipping network from 2017 to 2019.
Graph Metrics Computation
We analyzed the shipping network by calculating various graph metrics for each node. These metrics included betweenness centrality, closeness centrality, and PageRank. The detailed equations for calculating these metrics are listed in Equations 1416, and 17 in the Methods. In this paper, betweenness centrality quantifies the frequency with which a port serves as an intermediary on the shortest paths between other ports — ports with high betweenness centrality play a critical role as bridges within the shipping network. Additionally, closeness centrality calculates the inverse of the sum of the shortest distances from a node to all other nodes — ports with a high degree of closeness are easily accessible from all other ports in the shipping network. Furthermore, PageRank [47] identifies essential nodes in a graph — ports with high metric values are more influential and likely to be frequently visited by ships from other important ports.
Port Name $C_{B}(i)$ $C_{C}(i)$ $P_{ij}$ Water Body Country Region
As Suways 0.018 0.532 0.003 Gulf of Suez Egypt Northern Africa
Link Prediction in the Shipping Network
Subsequently, during the shipping network analysis, we observed the presence of disconnected components. These disconnected segments pose a challenge for our proposed framework, which relies on integrating features across all possible destinations from each source port. Such disconnections can compromise the model’s ability to generate accurate or well-defined predictions for shipping flows between isolated areas. Thus, we transformed the original network into a complete graph, denoted as $G^{\prime}$, and assigned a small value to weigh the new links. This action mitigates the issues arising from data sparsity and establishes a uniform data structure, thereby enhancing the robustness of the flow estimation framework. With the fully connected shipping network $G^{\prime}$, we proceed with the flow estimation framework by forecasting whether a trajectory exists by solving a link prediction problem in $G^{\prime}$, which can be framed as a binary classification task within machine learning models. This action gives concrete source-destination pairs well-prepared for building feature sets, modeling the gravity structure, and introducing the deep learning framework.
To perform link prediction, we first separated the shipping data from 2019 for testing and retained the data from 2017 and 2018 for training and validation. Then, we prepared features for the classification task. We calculated the Haversine distances (see Methods) between every pair of ports. However, Haversine distances only provide the geodesic approximation and cannot capture the real sea routes that the ships have traveled. Therefore, we also computed the sea-route distances between port pairs and obtained a more accurate representation using a Python package that models the shortest routes and calculates the sea route distances using historical trajectories [48]. Finally, we used these distances to evaluate the importance level $I_{ij}$ for each edge $\langle i,j\rangle$ in the complete graph $G^{\prime}$. Inspired by straightness centrality measuring the node connectivity by the straightness of the shortest distance [49], this metric combines the normalized flow size $\hat{w}_{ij}$ and the normalized Haversine distance $\hat{d_{ij}^{E}}$, using a small constant $\epsilon$ to prevent division by zero. Connections with more shipping flows and shorter distances are considered more important:
Next, we incorporated the Haversine distance, sea route distance, and edge importance into the feature set to determine the existence of a trajectory between two ports. Upon labeling the real (true class) and pseudo (false class) links in the complete network $G^{\prime}$, we observed a significant class imbalance (2.3% real links and 97.7% pseudo links); consequently, we employed stratified sampling of the pseudo links based on their spatial distribution to balance the number of examples in each class in a manner that preserved the data’s characteristics. Machine learning models were then engaged to perform binary classification, utilizing 75% of data from 2017 and 2018 for training and 25% for validation. During the training phase, a 5-fold grid search cross-validation was implemented to fine-tune the hyperparameters of each model. Finally, we tested the models on unseen data from 2019, reinforcing the validity of our approach.
@@
Figure 4: validation and test accuracy of classifiers in the trajectory link prediction task. (a) Performance of the classification task incorporates Haversine distance, sea route distance, and edge importance as features. (b) Performance of classification task without the edge importance features.
@@
Figure 3(a) displays the validation and test accuracy across different classification models. It can be observed that apart from Logistic Regression at 89%, all other models achieved high accuracy (above 98%) on both the validation and the test datasets. We further investigated the discrepancies between the model with the highest performance (Random Forest at 100%) and the model with the lowest accuracy (Logistic Regression at 89%). The result revealed that the links primarily missed by Logistic Regression were those with only 1 or 2 trips and, to a lesser extent, 3-trip links. We also observed that the high accuracy of some models could be attributed to the inclusion of normalized edge weights $w_{ij}$ in $G^{\prime}$ (i.e., the normalized number of trips) in the edge importance calculation as per Equation 1. Consequently, we ran predictions without edge importance as a feature to mitigate this effect for comparison. Figure 3(b) presents the validation and test accuracy when relying solely on Haversine and sea route distance as features. The accuracy of these models ranges from 75% to 79%. Notably, the models’ performance decreases by 16% to 20% compared to models that use edge importance as a feature. Retaining critical details from trade fluxes directly proportional to the link’s existence is the key function of the edge importance feature. However, we have observed that more than the edge importance is needed to create a self-contained inference model. The challenge lies in selecting a model that can utilize the extra information without overfitting the data.

### v3

# Enhancing Global Maritime Traffic Network Forecasting with Gravity-Inspired Deep Learning Models

@@ Abstract
Aquatic non-indigenous species (NIS) pose significant threats to biodiversity, disrupting ecosystems and inflicting substantial economic damages across agriculture, forestry, and fisheries. Due to the fast growth of global trade and transportation networks, NIS has been introduced and spread unintentionally in new environments. This study develops a new physics-informed model to forecast maritime shipping traffic between port regions worldwide. The predicted information provided by these models, in turn, is used as input for risk assessment of NIS spread through transportation networks to evaluate the capability of our solution. Inspired by the gravity model for international trades, our model considers various factors that influence the likelihood and impact of vessel activities, such as shipping flux density, distance between ports, trade flow, and centrality measures of transportation hubs. Accordingly, this paper introduces transformers to gravity models to rebuild the short- and long-term dependencies that make the risk analysis feasible. Thus, we introduce a physics-inspired framework that achieves an 89% binary accuracy for existing and non-existing trajectories and an 84.8% accuracy for the number of vessels flowing between key port areas, representing more than 10% improvement over the traditional deep-gravity model. Along these lines, this research contributes to a better understanding of NIS risk assessment. It allows policymakers, conservationists, and stakeholders to prioritize management actions by identifying high-risk invasion pathways. Besides, our model is versatile and can include new data sources, making it suitable for assessing international vessel traffic flow in a changing global landscape.
@@ Introduction
Globalization has rapidly increased marine shipping activities in the last decades. According to a statistics report, container shipping has increased by 24 times in tonnage from 1980 to 2020 [1]. This significant growth has not only boosted global trade but has also raised environmental concerns, including increased ocean pollution and the introduction of NIS through ballast water across marine ecosystems [2]. Recent research has revealed a strong correlation between the introduction of NIS and the movement of ships through shipping networks. These studies [3, 4] have utilized data from the Automatic Identification System (AIS), a location tracking system on ships that allows them to share their positions in real-time [5]. This technology allows researchers to track individual/collective ships [6, 7, 8, 9], predict small- or larger-scale shipping activities [10, 11], and assess the risk of introducing NIS through ballast water [3] discharge.
@@
AIS data has emerged as a promising source of information for studying maritime ship traffic patterns. The mobility intrinsic in AIS data associated it with studies of mobility flows by using Origin-Destination (OD) models that combine physics with statistical mechanics, including the classical gravity model, which, inspired by Newton’s Law of Universal Gravitation, measures the attractive force between two objects based on their masses and the distance between them [12]. The gravity-inspired OD models were introduced in early human mobility and migration studies [13, 14], where the number of commuters between two locations relies on population size and distances between origins and destinations as features.
The gravity theory describes the interaction between two entities proportionally to their masses and inversely proportional to their distance [12]. It permeates many areas of study that go beyond mobility and migration, such as the spreading of epidemics [15, 16], commercial trading [17], communication network [18] and cargo shipping networks [4, 19, 20] modeling. Although the gravity model has been widely used to model real-world problems, recent studies have shown that it may not be sufficient for capturing complex patterns in various scenarios [21]. Relying solely on mass and distance as the critical factors of the model could lead to failures in accurately representing patterns [22]. Also, the lack of limits on the flow size can lead to the predicted flow size larger than the source “masses”, making the model predictions unreasonable and challenging to interpret [23, 24]. Additionally, it is difficult to deal with multiple adjustable parameters in the model for the prediction without enough previously observed data. Nonetheless, the gravity model has been prevalent for many years and remains a popular tool for modeling various phenomena.
Beyond gravity models, radiation absorption is another physics-inspired OD model to study mobility patterns [23, 25, 26]. Unlike gravity models, radiation models are based on principles seen in radiation and absorption processes from physics. The radiation model overcomes several limitations of gravity models in predicting mobility flows. While gravity models contain adjustable variables that may be difficult to define, the radiation-absorption model simplifies this by emphasizing distance as the primary feature while considering the population density and offers a parameter-free approach that resolves the problem of having multiple parameters in the gravity model. Besides, it limits the number of flows by introducing the total number of individuals departing from the source location [23, 26]. Further physics-inspired studies used field theory for abstracting vector fields of daily commuting flows [27], while others translated field theory-based mobility to deep learning models for achieving better interpretation of spatiotemporal features in mobility patterns [28]. Moreover, different studies have incorporated multiple relative factors, such as employment and urban infrastructure, that capture more important characteristics to improve prediction accuracy and adaptability of the model [24, 29]. The evolving nature of artificial intelligence and the rise of large language models have brought about new technologies. Therefore, improving the state of the art by using more capable technologies over new arrangements and combinations of data is essential. This becomes more evident when considering evolving factors such as climate change and recurrent anthropogenic effects observed on the oceans.
@@
Deep learning techniques have become increasingly popular in various applications due to their ability to recognize patterns by fine-tuning multiple parameters. These techniques have been used to forecast vessel trajectories in the ocean, predict patient trajectories in hospitals, track the spread of epidemics, and many other applications [30, 31, 32]. Understandably, coupling machine or deep learning capabilities for pattern recognition with physics-inspired OD models can imply a higher proficiency in capturing and predicting complex scenarios. For instance, a study in human mobility used deep learning methods with the gravity model [29]. The resulting composite model, called the Deep Gravity Model, has expanded the standard feature set of conventional gravity models, which typically incorporated population size and distance between OD locations, to include a variety of parameters characterizing the origins and destinations such as land-use patterns and the presence of retail and healthcare amenities. However, the simplicity of the multilayer perceptrons (MLPs) structure in Deep Gravity models presents certain limitations. Specifically, accurately capturing the complex multivariate relationships inherent in mobility flows can be challenging due to the composition of multiple functions, which provides an opportunity for further research efforts.
@@
Along these lines, this paper focuses on predicting the OD fluxes of marine vessels to gain insight into global shipping behavior and revert such information to the risk assessment of NIS cases through BWRA for performance comparison and evaluation against state-of-the-art machine-learning-driven solutions. Ballast water is essential in maritime operations, particularly for container ships. As illustrated in Figure 1, while ballast water helps stabilize the ship’s load, local species can enter the ballast tanks and travel long distances when the water is taken in and discharged. Therefore, forecasting shipping patterns is critical to understanding the risk of spreading NIS. To this end, we propose a gravity-informed model where the shipping fluxes are considered as “mass”, and the vessel traffic flows are inversely proportional to the distances traveled by ships. We have enriched our model with relevant features from shipping activities, including bilateral trade data between countries [33] and graph metrics extracted from the global shipping network. Our deep learning model, known as the Transformer Gravity, relies on the transformer architecture [34] and is proficient in capturing local and global data dependencies through self-attention mechanisms. This mechanism enables the model to weigh the importance of different parts of the sequence of input data (i.e., information about maritime vessel movements), assigning varying degrees of attention depending on the relevance of each input part while generating the output. As a result, the model we propose can better discern and incorporate short- and long-term dependencies in vessel traffic flows, making it more sensitive to the complex and dynamic patterns in maritime vessel movements.
@@
As part of our proposed framework, we have employed a machine learning classifier that proceeds the flow estimation process by identifying the most probable OD pairs for gravity models. In this step, we first construct a fully connected shipping network to fill gaps in the original network. We weigh the edges based on distance and the low likelihood of a vessel traveling between the two locations (see Methods). The classifier then evaluates each link, learning how to filter out those less likely to exist based on critical attributes, including distance and edge importance (see Methods). This allows only highly probable flows, based on prior knowledge, to be fed into the gravity-based models, where the final flow estimations take place with the aid of the gravity-based model. In this sense, we have conducted experiments using the (a) Transformer Gravity, (b) Deep Gravity, and (c) shallower-layered variants of Deep Gravity Models. We utilized regression models in machine learning for performance comparison. Our results demonstrate that the Transformer Gravity model significantly outperforms all the other approaches as it achieves an average Common Part of Commuters (CPC) of 86.4%, representing an improvement higher than 10% in the model output certainty in contrast to the benchmark Deep Gravity and its shallower-layered variants, and nearly 50% improvement compared to machine learning regression models. The results we have obtained are not only due to the proposed model but also to the pipeline outputs of the most probable connections that provide prior knowledge to the gravity-informed models (see Methods). We have significantly improved performance by incorporating prior knowledge about potential destinations and using the attention mechanism and the traditional gravity-informed model for mobility flow estimation.
@@ Results
We start this study by constructing a directed and weighted international shipping network based on global port visits data from 2017 to 2019, which was derived from AIS data as processed in the study of Carlini et al. [35]. Ports and shipping connections were represented as nodes and edges in the shipping network, and the World Port Index (WPI) [36] was used for port identification. Figure 2 provides an overview of global shipping connections in the three years from 2017 to 2019, which shows the distribution of a total of $2,304$ ports (i.e., nodes), $141,179$ shipping connections (i.e., edges) and $3,573,979$ trips participating in shipping activities. The number of trips along the same shipping connection was used as the edge weight.
Figure 2: Global Shipping Network 2017–2019. Edge color and thickness are relative to the number of shipping activities per route, with darker blue and bolder lines indicating routes with higher activity levels. Port sizes in color-coded circles are scaled according to shipping fluxes to highlight their port capacity.
Figure 3: The experimental pipeline for predicting ship traffic flows with gravity-informed models, including the assessment of environmental similarity used on the ballast water risk assessment case study. The process begins with identifying and verifying predicted links between source ports and target regions. Subsequently, key features are collected, and graph metrics are extracted so they can be used to inform the gravity-based model used to forecast ship traffic flow sizes. Concurrently, environmental data pertinent to the study regions is gathered and evaluated with the aim of computing environmental similarity metrics.
@@
A detailed description of the methodological steps for creating the shipping networks, preparing the data for link prediction tasks, feature extraction, analysis, and augmentation is included and extensively discussed in the Methods section at the end of this paper. It is noteworthy that our study was conducted during significant events (e.g., war, global pandemic, and economic sanctions) that could have influenced the underlying shipping data. However, all results we discuss use the same methodology, ensuring that any external factors would have had a homogeneous impact on the comparative performance of the models.
Network Analysis and Link Prediction
We conducted an analysis of the shipping network to extract graph metrics. These metrics were used as part of the feature set for gravity-informed predictive models, which helped forecast traffic flow sizes on the shipping network. We implemented link prediction as a binary classification task within the fully connected shipping network to identify the most probable shipping connections and mitigate the effect of missing links on the gravity model. These predictions are important for informing the model about the specific origin-destination (OD) pairs involved, as shown in the first two stages in Figure 3.
@@
Figure 4: Validation and test accuracy of classifiers in the trajectory link prediction task: (a) The performance of the classification task includes Haversine distance, sea route distance, and edge importance as features; and, (b) The classification task is carried out without the edge importance features.
Figure 5: Framework of the Transformer Gravity model. The process starts with input sequences that are embedded and passed through self-attention blocks with multi-head attention, dropout, and layer normalization. This is followed by feed-forward blocks containing linear layers and dropout, resulting in the output sequence. The cross-entropy loss with log-softmax is used for training, while the Common Part of Commuters (CPC) is used for evaluation by incorporating commuting patterns $O_{i}$ from the input data.
@@
Figure 4(a) displays the validation and test accuracy across different classification models for the binary classification-based link prediction task. It can be observed that apart from Logistic Regression at 89%, all other models achieved high accuracy (above 98%) on both the validation and the test datasets. We further investigated the discrepancies between the model with the highest performance — Random Forest at 100% — and the model with the lowest accuracy — Logistic Regression at 89%. The result revealed that the links primarily missed by Logistic Regression were those with only 1 or 2 trips and, to a lesser extent, 3-trip links. We also observed that the high accuracy of some models could be attributed to the inclusion of normalized edge weights $w_{ij}$ (i.e., the normalized number of trips) in the fully connected shipping network $G^{\prime}$ in the edge importance calculation as per Equation 8. Consequently, we ran predictions without edge importance as a feature to mitigate this effect for comparison. Figure 4(b) presents the validation and test accuracy when relying solely on Haversine and sea route distance as features. The accuracy of these models ranges from 75% to 79%. Notably, the models’ performance decreases by 16% to 20% compared to models that use edge importance. Retaining critical details from fluxes directly proportional to the link’s existence is the key function of the edge importance feature, but creating a self-contained inference model involves more than just edge importance. The challenge lies in selecting a model that can incorporate that information without overfitting the data.